OpenStudy (anonymous):

what is the circumference of a circle with a radius equaling 5 centimeters in length

OpenStudy (anonymous):

Circumference of a circle=\(2\pi r\) Can you do nw?

OpenStudy (anonymous):

\[\large Circumference = 2 \pi r\] Here , r = 5 and \(\pi\) = 3.14 \[\large \color{green}{C = 2 \times 3. 14 \times 5 = ?}\]